When an investor evaluates a 6-month compound interest of 1.500.000 TL with an interest rate of 15 percent, how much will it be with interest after 4 years?

 


1
Expert's answer
2021-01-20T13:52:54-0500

Amount"=1,500,000"

semi-annual rate "=15/2=7.5\%"

number of periods "=4\times 2=8"

values in 4 years = "=1,500,000\times (1+0.075)^{8}=2675216.74"

amount of interest "=2675216-1500000=1,175,216.74"